# RL in Latent MDPs is Tractable: Online Guarantees via Off-Policy Evaluation

### v1

@@ 7 Conclusion and Future Work
In this work, we presented the first sample-efficient algorithm for LMDPs, resolving an open question of efficient exploration with latent contexts. While our result is specialized, we believe our new perspectives and techniques on deriving online guarantees through the lens of OPE can be useful for a broader range of interactive learning, and, especially, partially observed problems. While resolving the open problem, there are few remaining questions for the LMDP setting.
@@
The OPE lemma derived in this work (Lemma 4.2) assumes the behavior policy is a segmented policy with intervention at different checkpoints. While this result allows us to provide guarantees on LMDP-OMLE and prove it learns a near-optimal policy, this result is restrictive. Specifically, it does not provide general guarantees for OPE nor makes it possible to derive regret guarantees. Can we evaluate $\pi\in\Pi$ without policy-switching or intervention when the behavioral policy is a generic history-dependent policy $\psi\in\Pi$? Further, is there an algorithm with provable $\mathrm{poly}(S,A)^{M}\cdot\sqrt{T}$ regret for the general LMDP setting?
@@
Our result gives a worst-case guarantee. Yet, practical instances may be much simpler under different set of assumptions (e.g., [43]). Is there an alternative notion of separation between contexts that may break the curse of contexts $M$? e.g., coverage-coefficient with a shorter length $d=o(M)$ of state-action tuples might be sufficient to decode the latent context. Deriving new conditions can be of great importance for real-world applications. Further, developing practical RL methodologies for the LMDP setting remains an unexplored challenge with significant importance for numerous applications. These are remained to be explored in future works.
@@ Appendix C Proofs for Section 4
We continue from the conclusion of Lemma 5.1, and the remaining step is to ensure that LMDP-OMLE terminates after $K=MS^{2}H\log(1/\gamma)$ iterations where $\gamma=\epsilon_{\texttt{test}}^{2}/(H^{2d})$ without the initial coverage assumption (8). We consider a perturbed model $\hat{\theta}^{*}=(\{w_{m},\hat{T}_{m},R_{m}\}_{m=1}^{M})$ where
@@
$\texttt{TV}(\mathds{P}_{\hat{\theta}^{*}}^{\pi},\mathds{P}_{\theta^{*}}^{\pi})(\mathcal{T})$ $\leq\textstyle\sum_{m}w_{m}\cdot\texttt{TV}(\mathds{P}_{\hat{\theta^{*}}}^{\pi},\mathds{P}_{\theta^{*}}^{\pi})(\mathcal{T}|m)\leq 2S\gamma H.$

### v2

@@ 7 Conclusion and Future Work
In this work, we presented the first sample-efficient algorithm for LMDPs, resolving an open question of efficient exploration with latent contexts. While our result is specialized to LMDPs, we believe our new perspectives and techniques on deriving online guarantees through the lens of OPE can be useful for a broader range of interactive learning, and, especially, partially observed problems. While resolving the open problem, there are a few remaining questions for the LMDP setting.
@@
The OPE lemma derived in this work (Lemma 4.2) assumes the behavior policy is a segmented policy with intervention at different checkpoints. While this result allows us to provide guarantees on LMDP-OMLE and prove it learns a near-optimal policy, this result is restrictive, in that it does not provide general guarantees for OPE nor makes it possible to derive regret guarantees. In particular, can we evaluate $\pi\in\Pi$ without policy-switching or intervention when the behavioral policy is a generic history-dependent policy $\psi\in\Pi$? Further, is there an algorithm with provable $\mathrm{poly}(S,A)^{M}\cdot\sqrt{T}$ regret for the general LMDP setting?
@@
Our result gives a worst-case guarantee. Yet, practical instances may be much simpler under different set of assumptions e.g., with provided side-information [66, 44] or additional structural assumptions [41, 63, 14]. Deriving new conditions can be of great importance for real-world applications, e.g., there could be more practical notion of separation, or the set of instances that allows the notion of coverage-coefficient with a (significantly) shorter length $d=o(M)$ of state-action tuples. Further, developing practical RL methodologies for the LMDP setting remains an unexplored challenge with significant importance for numerous applications. These are remained to be explored in future works.
@@ Appendix C Proofs for Section 4
We continue from the conclusion of Lemma 5.1, and the remaining step is to ensure that LMDP-OMLE terminates after $K=MS^{2}H\log(1/\gamma)$ iterations where $\gamma=\epsilon_{\texttt{test}}^{2}/(H^{2d})$ without the initial coverage assumption (8). We consider a perturbed model $\hat{\theta}^{*}=(\{w_{m}^{*},\hat{T}_{m},R_{m}^{*}\}_{m=1}^{M})$ where
@@
$\texttt{TV}(\mathds{P}_{\hat{\theta}^{*}}^{\pi},\mathds{P}_{\theta^{*}}^{\pi})(\mathcal{T})$ $\leq\textstyle\sum_{m}w_{m}^{*}\cdot\texttt{TV}(\mathds{P}_{\hat{\theta^{*}}}^{\pi},\mathds{P}_{\theta^{*}}^{\pi})(\mathcal{T}|m)\leq 2S\gamma H.$